One of my questions I submitted earlier was sent in error, can you still help me with this last question?
Thanks

Boys and girls. Alice and Bill are planning to have three children. (Assume it is equally likely for a boy or a girl to be born.) Give the probability as a decimal.
(Fill in the blank below and give your answer to 3 decimal places.)
What is the probability that all three of their children will be girls?
The probability that all three children will be girls is .


Found 3 solutions by jim_thompson5910, stanbon, ikleyn:
Answer by jim_thompson5910(35256)

B = child is a boy
G = child is a girl

P(B) = 1/2 = 0.5
P(G) = 1/2 = 0.5

Because the events are independent, ie the gender of child1 doesn't affect the gender of child2 or child3, this means

P(3 girls) = P(girl and girl and girl)
P(3 girls) = P(girl)*P(girl)*P(girl)
P(3 girls) = P(G)*P(G)*P(G)
P(3 girls) = 0.5*0.5*0.5
P(3 girls) = 0.125

Answer by stanbon(75887)

Boys and girls. Alice and Bill are planning to have three children. (Assume it is equally likely for a boy or a girl to be born.) Give the probability as a decimal.
P(boy) = 1/2 = 0.5 ; P(girl) = 0.5
--------------------------------------------
(Fill in the blank below and give your answer to 3 decimal places.)
What is the probability that all three of their children will be girls?
The probability that all three children will be girls is 0.5^3 = 0.625
----------------
Cheers,
Stan H.
----------

Answer by ikleyn(52800)